OpenStudy (anonymous):

Find the points of the curve at which the tangent line to the function y=(cosx)/(2+sinx) is horizontal. I think i've mostly figured the problem out.

OpenStudy (anonymous):

Ok i know the derivative equal to 0 will give me my answer if i solve for x. So first I did quotient rule

OpenStudy (anonymous):

and got \[\frac{ (-sinx)(2+sinx)-(cosx)(cosx) }{ (2+sinx)^2 }\]

OpenStudy (anonymous):

I then simplified and got \[-2sinx-\sin^2x-\cos^2x=0\]

OpenStudy (anonymous):

I just don't know how to solve for x from this